# Awakening Augmented Generation: Learning to Awaken Internal Knowledge of Large Language Models for Question Answering

### v4

@@ Abstract
Huanxuan Liao1,2, Shizhu He1,2, Yao Xu1,2, Yuanzhe Zhang1,2, Kang Liu1,2, Shengping Liu3, Jun Zhao1,2 1 The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3 Unisound, Beijing, China liaohuanxuan2023@ia.ac.cn {yao.xu, shizhu.he, kliu, jzhao}@nlpr.ia.ac.cn
@@ 3.3 Implicit Awakening with Hypernetwork
The hypernetwork architecture for generating LoRA weights is detailed in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank projections of layer $k$ associated with the $\mathcal{Q}$, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the $\mathcal{V}$. The hypernetwork, denoted as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vector obtained using the model’s encoder and reduced in dimensionality via a whitening algorithm Su (2021). The term $idx_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding, differentiating between layers and $\mathcal{QV}$. Each hypernetwork is characterized by weights $W_{d}$ and $W_{u}$, representing the down and up projections, respectively. The hypernetwork equations for $D^{\{q,v\}}$ is expressed as follows:

### v5

@@ Abstract
Huanxuan Liao1,2, Shizhu He1,2thanks: Corresponding author, Yao Xu1,2, Yuanzhe Zhang1, Shengping Liu3, Kang Liu1,2, Jun Zhao1,2 1 The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3 Unisound, Beijing, China liaohuanxuan2023@ia.ac.cn {yao.xu, shizhu.he, kliu, jzhao}@nlpr.ia.ac.cn
@@ 3.3 Implicit Awakening with Hypernetwork
The hypernetwork architecture for generating LoRA weights is detailed in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank down and up projections of layer $k$ associated with the Query matrix $W_{\mathcal{Q}}$ in the attention module, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the Value matrix $W_{\mathcal{V}}$. The hypernetwork, denoted as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vector obtained using the model’s encoder and reduced in dimensionality via a whitening algorithm Su (2021). To achieve this whitening transformation, we first compute the mean of the vector $\mu=\frac{1}{N}\sum_{i=1}^{N}x_{i}$ and center the data by subtracting $\mu$ from each vector $x_{i}$. Next, we calculate the covariance matrix $C$ of the centered vectors $\tilde{x}_{i}=x_{i}-\mu$, which is given by $C=\frac{1}{N}\sum_{i=1}^{N}\tilde{x}_{i}\tilde{x}_{i}^{T}$. We then perform Singular Value Decomposition (SVD) on the covariance matrix: $C=U\Lambda U^{T}$, where $U$ contains the eigenvectors and $\Lambda$ is a diagonal matrix of eigenvalues. The transformation matrix $W$ is derived from the eigenvalue decomposition as $W=U\Lambda^{-1/2}$, where $\Lambda^{-1/2}$ scales the eigenvectors by the inverse square root of their corresponding eigenvalues. Thus, applying the transformation $\tilde{x}_{i}=(\tilde{x}_{i})W$ not only centers the data around zero but also results in a covariance matrix that is equivalent to the identity matrix, ensuring that the transformed vectors are uncorrelated and have unit variance. The term $idx_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding, differentiating between layers and $\mathcal{QV}$. Each hypernetwork is characterized by weights $W_{d}$ and $W_{u}$, representing the down and up projections, respectively. The hypernetwork equations for $D^{\{q,v\}}$ is expressed as follows:
@@ 5 Conclusion
Acknowledgements
This work was supported by National Key R&D Program of China (No. 2022YFF0711900) and the National Natural Science Foundation of China (No.62376270, No.62276264). This work was supported by the Youth Innovation Promotion Association CAS.